# Poutine: Vision-Language-Trajectory Pre-Training and Reinforcement Learning Post-Training Enable Robust End-to-End Autonomous Driving

### v3

@@ 1 Introduction
Vision–language models (VLMs) have emerged as a powerful means of coupling visual perception with the world knowledge and common sense reasoning acquired from internet-scale pre-training [3, 1, 7, 10, 9]. For autonomous vehicles, such multimodal reasoning is most valuable in long-tail situations—rare but safety-critical events that dominate operational risk and have limited coverage in conventional driving corpora. Nevertheless, the empirical study of VLMs for driving has so far been restricted largely to nominal driving benchmarks such as nuScenes [6], where high-level semantic reasoning is seldom required and the benefits of language grounding remain unclear [12, 15, 24, 30, 22, 5, 32, 31, 37, 13, 34, 41, 20, 14, 38]. The Waymo Vision-Based End-to-End Driving (WOD-E2E) dataset [36], along with its accompanying challenge, presents a new opportunity for researchers to test trajectory planning in more challenging curated long-tail scenarios that account for less than 0.003% of daily driving. This is in stark contrast to the relatively boring scenarios present in other open-loop planning benchmarks. Hence, the WOD-E2E dataset provides a more suitable testbed for investigating whether the generalist knowledge embedded in VLMs can translate into safer and more reliable driving policies.
To address this question, we introduce Poutine [^2] Named after the Québécois dish which combines fries, gravy, and cheese curds., a 3B-parameter VLM for end-to-end driving, trained via a simple two-stage pipeline that fine-tunes a pre-trained Qwen2.5-VL model [4]. In Stage 1, Poutine-Base learns base driving capabilities by performing standard next-token prediction over vision, language, and trajectory (VLT) input tokens to predict both a driving explanation and future trajectory, where the trajectories are represented as text. The training corpus comprises 83 hours of nominal Japanese driving from the public CoVLA dataset [2] and 11 hours of long-tail driving from WOD-E2E. All language annotations are generated automatically by a 72B-parameter VLM, removing the need for manual labeling and yielding a fully self-supervised pre-training procedure that is straightforward to scale. In Stage 2, we refine Poutine-Base with Group Relative Policy Optimization (GRPO) using less than 500 human preference-labeled frames from the WOD-E2E validation set. Despite the modest supervision budget, this lightweight reinforcement-learning (RL) step materially improves policy performance in the long-tail.
@@
Contributions: (1) We present Poutine, a simple and scalable VLM-based E2E driving policy that combines large-scale VLT pre-training with lightweight GRPO fine-tuning. We show that the combination of next-token prediction pre-training followed by RL finetuning that has proven successful in other domains [23, 11] can yield robust E2E motion planning policies. (2) We show that Poutine-Base enables zero-shot transfer to unseen geographic regions, as the model variant pre-trained exclusively on Japanese data generalizes zero-shot to the US despite the substantial distribution shift (i.e., left- versus right-hand driving). (3) Poutine secures first place in the 2025 Waymo Vision-Based End-to-End Driving Challenge, establishing a new benchmark for long-tail autonomy.
@@ 2 Related Work
VLMs for End-to-End Driving. Several prior works finetune VLMs for end-to-end driving in the open-loop [30, 14, 34, 31, 5, 13, 41, 39] and closed-loop [28, 27, 40, 35, 8, 26] setting. In both settings, evaluations are performed on nominal benchmarks such as nuScenes or on synthetic driving benchmarks like CARLA whose scenario diversity is limited; none report results on a curated real-world long-tail corpus comparable to WOD-E2E, which is the focus of the present work. A separate line of research employs VLMs as auxiliary components—either as natural-language conditioners for conventional end-to-end stacks [15, 32] or as teachers whose features are distilled into end-to-end stacks [38, 12, 24]. Poutine is a VLM finetuned directly for end-to-end driving, but unlike prior work, we additionally leverage RL post-training to enhance the model’s driving capabilities in long-tail scenarios.
@@
Prior works use custom perception backbones [34, 31, 41] and action headers [8, 27, 28] that are aligned to language space. To maintain the simplest design, we finetune an off-the-shelf VLM without any architectural modifications —leveraging both its vision encoder and language decoder—and predict the trajectory directly as natural language tokens. This alleviates the need for custom visual perception backbones or trajectory headers. The most comparable prior work is EMMA [14], which likewise fine-tunes an off-the-shelf VLM for end-to-end driving. In contrast to EMMA’s purely supervised fine-tuning approach, Poutine supplements supervised learning with preference-based RL, yielding improved driving capabilities in long-tail scenarios.
@@
RL Post-Training for End-to-End Driving. TrajHF [17] employs preference-based reinforcement learning to align open-loop trajectory predictions with human preferences, but its fine-tuning stage consumes 78 000 labelled frames and is applied to a model that lacks large-scale VLM pre-training. In contrast, Poutine bootstraps from both internet-scale VLM pre-training and our vision–language–trajectory (VLT) pre-training, achieving more human-aligned planned trajectories with only 479 preference-labelled frames. Group Relative Policy Optimization (GRPO) has recently proved effective for boosting reasoning capabilities in language models [11] and has been adopted by AlphaDrive [16] to refine high-level natural-language behaviour planners. AlphaDrive, however, does not generate planned trajectories. By contrast, Poutine leverages GRPO to directly improve trajectory planning performance.
@@
Concurrent to our work are AutoVLA [42] and Drive-R1 [18], which also explore combining self-supervised pretraining with GRPO fine-tuning for VLM-based trajectory planning in autonomous driving. However, AutoVLA reports substantially lower performance on the WOD-E2E benchmark, suggesting that its physical action token generation component may negatively impact overall performance. Meanwhile, Drive-R1 is only evaluated on nuScenes, a benchmark that consists primarily of nominal driving scenarios.
@@
Figure 1: (Top row) The Poutine model adopts an off-the-shelf 3B-parameter VLM with no additional components, and simply encodes trajectories in text space. (Bottom row) The proposed pipeline to train an end-to-end planner for robust autonomous driving. Datasets shown in the figure (i.e., CoVLA [2] and WOD-E2E [36]) indicate those used in our experiments. Fine-tuning of the Poutine model follows the highlighted stages with the vision encoder kept frozen.
@@ 3 Poutine
Figure 1 provides an overview of our trajectory‑planning framework for autonomous driving. Our model builds on an autoregressive, transformer‑based VLM pre‑trained on internet‑scale text and image data. In this work, we adopt the Qwen2.5-VL 3B Instruct model [4], which combines a Vision Transformer (ViT) encoder with a decoder‑only transformer language model. No additional architectural components or custom tokenizers are introduced to the Qwen model. As such, we treat trajectory planning as a next‑token prediction task over the original token vocabulary, thus representing trajectories directly as text.
@@ 3.3 Reinforcement Learning Post-Training
We introduce a reinforcement learning post-training stage leveraging the GRPO algorithm [29] to fine-tune the Poutine-Base model on a set of high quality examples. In this stage, we fine-tune only on the future trajectory prediction task and omit outputting a structured CoT reasoning trace. This design decision significantly improves inference latency without impacting performance, as will be made clear in Section 4.
@@ 4 Results
Swin-Trajectory [25] 7.54 6.68 2.81 1.21
AutoVLA [42] 7.56 6.94 2.96 1.35
UniPlan [19] (DiffusionDrive) 7.69 6.65 2.99 1.31
DiffusionLTF [21] 7.72 6.41 2.89 1.36
HMVLM [33] 7.74 6.73 3.07 1.33
@@
Datasets. For VLT pre-training, we leveraged the CoVLA [2] and WOD-E2E datasets. CoVLA contains 10,000 front-view, 30 s driving videos recorded in Japan at 20 Hz with corresponding ego trajectory information. For each frame, we stored the future 5 s ego trajectory subsampled to 4 Hz to be consistent with the WOD-E2E format. The WOD-E2E dataset provides 4021 long-tail driving scenarios of 20 s each, captured at 10 Hz from 8 multi-view cameras. Each scenario belongs to one of 10 (11 for test set) predefined long-tail scenario categories, such as pedestrians, intersections, cut-ins, and special vehicles, where the degree of ‘rarity’ significantly varies across scenarios. The official split allocates 2037 videos for training, 479 for validation, and 1505 for testing. Each frame contains a 4 s ego-vehicle past trajectory and a 5 s ground-truth future trajectory, both sampled at 4 Hz. Only the three front-facing camera images from the current frame in the WOD-E2E dataset were used as input to the model [^3] Although omitting side and rear views reduced contextual coverage, we found that the front images are sufficient for most scenarios and their smaller token count enabled substantially more optimization steps within our compute budget., whereas we used three consecutive frames from the preceding second for CoVLA (see Fig. 1 for an example). The raw images were downsampled to a maximum resolution of $H\times W=512\times 512$ pixels from their original sizes of $1208\times 1928$ (CoVLA) and $1079\times 972$ pixels (WOD-E2E). The WOD-E2E validation set contains 479 human preference-labeled frames, used as part of our RL fine-tuning stage.
@@
Note that while AutoVLA [42] shares significant overlap with Poutine, it demonstrates a limited improvement (+0.03 RFS) compared to Poutine’s substantial gain (+0.46 RFS) relative to the Waymo baseline. One possible explanation is that the driving-specific action tokenizer used by AutoVLA limits the ability to leverage world knowledge embedded in the base VLM, compared to directly encoding trajectories in text space.
Ablations. Several ablations were conducted to assess the various design decisions of Poutine using the WOD-E2E validation set. Table 3 analyzes pre-training ablations. VLMs that skip VLT pre-training (i.e., the two Qwen baselines) trail all Poutine variants, confirming the necessity of domain-specific pre-training. Training on CoVLA alone (Poutine-Base CoVLA) transfers zero-shot to Waymo with an RFS of 7.74, demonstrating strong cross-domain generalization. Removing CoVLA data (Poutine-Base No-CoVLA) or the auto-generated captions (Poutine-Base No Language) from training degrades performance to 7.95 and 7.94, respectively, showing that both large-scale pre-training and language supervision are beneficial. Consistent with prior works [14, 30], generating CoT at inference (Poutine-Base CoT) does not improve over the no-CoT Poutine-Base variant. Further investigation is required to determine the benefit of CoT for reasoning in long-tail scenarios at inference time. The final Poutine-Base model achieves an 8.12 RFS, nearly matching the 8.13 RFS attained with the expert WOD-E2E trajectories.

### v4

@@ 1 Introduction
Vision–language models (VLMs) have emerged as a powerful means of coupling visual perception with the world knowledge and common sense reasoning acquired from internet-scale pre-training [bai2023qwen, gpt42023openai, chen2024internvl, dubey2024llama3, reid2024gemini]. For autonomous vehicles, such multimodal reasoning is most valuable in long-tail situations—rare but safety-critical events that dominate operational risk and have limited coverage in conventional driving corpora. Nevertheless, the empirical study of VLMs for driving has so far been restricted largely to nominal driving benchmarks such as nuScenes [caesar2020nuscenes], where high-level semantic reasoning is seldom required and the benefits of language grounding remain unclear [hedge2025dima, jiang2024senna, pan2024vlp, sima2024drivelm, nie2024reason2drive, bai2024atlas, tian2024drivevlm, tian2024token, xing2025openemma, huang2024rdadriver, wang2024omnidrive, zhou2025opendrivevla, mao2023gptdriver, hwang2024emma, xu2024vlmad]. The Waymo Vision-Based End-to-End Driving (WOD-E2E) dataset [waymo2025e2e], along with its accompanying challenge, presents a new opportunity for researchers to test trajectory planning in more challenging curated long-tail scenarios that account for less than 0.003% of daily driving. This is in stark contrast to the relatively boring scenarios present in other open-loop planning benchmarks. Hence, the WOD-E2E dataset provides a more suitable testbed for investigating whether the generalist knowledge embedded in VLMs can translate into safer and more reliable driving policies.
To address this question, we introduce Poutine [^2] Named after the Québécois dish which combines fries, gravy, and cheese curds., a 3B-parameter VLM for end-to-end driving, trained via a simple two-stage pipeline that fine-tunes a pre-trained Qwen2.5-VL model [Qwen2.5-VL]. In Stage 1, Poutine-Base learns base driving capabilities by performing standard next-token prediction over vision, language, and trajectory (VLT) input tokens to predict both a driving explanation and future trajectory, where the trajectories are represented as text. The training corpus comprises 83 hours of nominal Japanese driving from the public CoVLA dataset [arai2025covla] and 11 hours of long-tail driving from WOD-E2E. All language annotations are generated automatically by a 72B-parameter VLM, removing the need for manual labeling and yielding a fully self-supervised pre-training procedure that is straightforward to scale. In Stage 2, we refine Poutine-Base with Group Relative Policy Optimization (GRPO) using less than 500 human preference-labeled frames from the WOD-E2E validation set. Despite the modest supervision budget, this lightweight reinforcement-learning (RL) step materially improves policy performance in the long-tail.
@@
Contributions: (1) We present Poutine, a simple and scalable VLM-based E2E driving policy that combines large-scale VLT pre-training with lightweight GRPO fine-tuning. We show that the combination of next-token prediction pre-training followed by RL finetuning that has proven successful in other domains [ouyang2022rlhf, deepseek2025r1] can yield robust E2E motion planning policies. (2) We show that Poutine-Base enables zero-shot transfer to unseen geographic regions, as the model variant pre-trained exclusively on Japanese data generalizes zero-shot to the US despite the substantial distribution shift (i.e., left- versus right-hand driving). (3) Poutine secures first place in the 2025 Waymo Vision-Based End-to-End Driving Challenge, establishing a new benchmark for long-tail autonomy.
@@ 2 Related Work
VLMs for End-to-End Driving. Several prior works finetune VLMs for end-to-end driving in the open-loop [sima2024drivelm, hwang2024emma, wang2024omnidrive, tian2024token, bai2024atlas, huang2024rdadriver, zhou2025opendrivevla, xu2024drivegpt4] and closed-loop [shao2024lmdrive, renz2025simlingo, zhang2024wisead, wang2023drivemlm, fu2025orion, paul2024legodrive] setting. In both settings, evaluations are performed on nominal benchmarks such as nuScenes or on synthetic driving benchmarks like CARLA whose scenario diversity is limited; none report results on a curated real-world long-tail corpus comparable to WOD-E2E, which is the focus of the present work. A separate line of research employs VLMs as auxiliary components—either as natural-language conditioners for conventional end-to-end stacks [jiang2024senna, tian2024drivevlm] or as teachers whose features are distilled into end-to-end stacks [xu2024vlmad, hedge2025dima, pan2024vlp]. Poutine is a VLM finetuned directly for end-to-end driving, but unlike prior work, we additionally leverage RL post-training to enhance the model’s driving capabilities in long-tail scenarios.
@@
Prior works use custom perception backbones [wang2024omnidrive, tian2024token, zhou2025opendrivevla] and action headers [fu2025orion, renz2025simlingo, shao2024lmdrive] that are aligned to language space. To maintain the simplest design, we finetune an off-the-shelf VLM without any architectural modifications —leveraging both its vision encoder and language decoder—and predict the trajectory directly as natural language tokens. This alleviates the need for custom visual perception backbones or trajectory headers. The most comparable prior work is EMMA [hwang2024emma], which likewise fine-tunes an off-the-shelf VLM for end-to-end driving. In contrast to EMMA’s purely supervised fine-tuning approach, Poutine supplements supervised learning with preference-based RL, yielding improved driving capabilities in long-tail scenarios.
@@
RL Post-Training for End-to-End Driving. TrajHF [li2025finetuning] employs preference-based reinforcement learning to align open-loop trajectory predictions with human preferences, but its fine-tuning stage consumes 78 000 labelled frames and is applied to a model that lacks large-scale VLM pre-training. In contrast, Poutine bootstraps from both internet-scale VLM pre-training and our vision–language–trajectory (VLT) pre-training, achieving more human-aligned planned trajectories with only 479 preference-labelled frames. Group Relative Policy Optimization (GRPO) has recently proved effective for boosting reasoning capabilities in language models [deepseek2025r1] and has been adopted by AlphaDrive [jiang2025alphadrive] to refine high-level natural-language behaviour planners. AlphaDrive, however, does not generate planned trajectories. By contrast, Poutine leverages GRPO to directly improve trajectory planning performance.
@@
Concurrent to our work are AutoVLA [zhou2025autovla] and Drive-R1 [li2025drive], which also explore combining self-supervised pretraining with GRPO fine-tuning for VLM-based trajectory planning in autonomous driving. However, AutoVLA reports substantially lower performance on the WOD-E2E benchmark, suggesting that its physical action token generation component may negatively impact overall performance. Meanwhile, Drive-R1 is only evaluated on nuScenes, a benchmark that consists primarily of nominal driving scenarios.
@@
Figure 1: (Top row) The Poutine model adopts an off-the-shelf 3B-parameter VLM with no additional components, and simply encodes trajectories in text space. (Bottom row) The proposed pipeline to train an end-to-end planner for robust autonomous driving. Datasets shown in the figure (i.e., CoVLA [arai2025covla] and WOD-E2E [waymo2025e2e]) indicate those used in our experiments. Fine-tuning of the Poutine model follows the highlighted stages with the vision encoder kept frozen.
@@ 3 Poutine
Figure 1 provides an overview of our trajectory‑planning framework for autonomous driving. Our model builds on an autoregressive, transformer‑based VLM pre‑trained on internet‑scale text and image data. In this work, we adopt the Qwen2.5-VL 3B Instruct model [Qwen2.5-VL], which combines a Vision Transformer (ViT) encoder with a decoder‑only transformer language model. No additional architectural components or custom tokenizers are introduced to the Qwen model. As such, we treat trajectory planning as a next‑token prediction task over the original token vocabulary, thus representing trajectories directly as text.
@@ 3.3 Reinforcement Learning Post-Training
We introduce a reinforcement learning post-training stage leveraging the GRPO algorithm [shao2024deepseekmath] to fine-tune the Poutine-Base model on a set of high quality examples. In this stage, we fine-tune only on the future trajectory prediction task and omit outputting a structured CoT reasoning trace. This design decision significantly improves inference latency without impacting performance, as will be made clear in Section 4.
@@ 4 Results
Swin-Trajectory [zhou2025swinfrajectory] 7.54 6.68 2.81 1.21
AutoVLA [zhou2025autovla] 7.56 6.94 2.96 1.35
UniPlan [liao2025diffusiondrive] (DiffusionDrive) 7.69 6.65 2.99 1.31
DiffusionLTF [Nguyen2025DiffusionLTF] 7.72 6.41 2.89 1.36
HMVLM [wang2025hmvlm] 7.74 6.73 3.07 1.33
@@
Datasets. For VLT pre-training, we leveraged the CoVLA [arai2025covla] and WOD-E2E datasets. CoVLA contains 10,000 front-view, 30 s driving videos recorded in Japan at 20 Hz with corresponding ego trajectory information. For each frame, we stored the future 5 s ego trajectory subsampled to 4 Hz to be consistent with the WOD-E2E format. The WOD-E2E dataset provides 4021 long-tail driving scenarios of 20 s each, captured at 10 Hz from 8 multi-view cameras. Each scenario belongs to one of 10 (11 for test set) predefined long-tail scenario categories, such as pedestrians, intersections, cut-ins, and special vehicles, where the degree of ‘rarity’ significantly varies across scenarios. The official split allocates 2037 videos for training, 479 for validation, and 1505 for testing. Each frame contains a 4 s ego-vehicle past trajectory and a 5 s ground-truth future trajectory, both sampled at 4 Hz. Only the three front-facing camera images from the current frame in the WOD-E2E dataset were used as input to the model [^3] Although omitting side and rear views reduced contextual coverage, we found that the front images are sufficient for most scenarios and their smaller token count enabled substantially more optimization steps within our compute budget., whereas we used three consecutive frames from the preceding second for CoVLA (see Fig. 1 for an example). The raw images were downsampled to a maximum resolution of $H\times W=512\times 512$ pixels from their original sizes of $1208\times 1928$ (CoVLA) and $1079\times 972$ pixels (WOD-E2E). The WOD-E2E validation set contains 479 human preference-labeled frames, used as part of our RL fine-tuning stage.
@@
Note that while AutoVLA [zhou2025autovla] shares significant overlap with Poutine, it demonstrates a limited improvement (+0.03 RFS) compared to Poutine’s substantial gain (+0.46 RFS) relative to the Waymo baseline. One possible explanation is that the driving-specific action tokenizer used by AutoVLA limits the ability to leverage world knowledge embedded in the base VLM, compared to directly encoding trajectories in text space.
Ablations. Several ablations were conducted to assess the various design decisions of Poutine using the WOD-E2E validation set. Table 3 analyzes pre-training ablations. VLMs that skip VLT pre-training (i.e., the two Qwen baselines) trail all Poutine variants, confirming the necessity of domain-specific pre-training. Training on CoVLA alone (Poutine-Base CoVLA) transfers zero-shot to Waymo with an RFS of 7.74, demonstrating strong cross-domain generalization. Removing CoVLA data (Poutine-Base No-CoVLA) or the auto-generated captions (Poutine-Base No Language) from training degrades performance to 7.95 and 7.94, respectively, showing that both large-scale pre-training and language supervision are beneficial. Consistent with prior works [hwang2024emma, sima2024drivelm], generating CoT at inference (Poutine-Base CoT) does not improve over the no-CoT Poutine-Base variant. Further investigation is required to determine the benefit of CoT for reasoning in long-tail scenarios at inference time. The final Poutine-Base model achieves an 8.12 RFS, nearly matching the 8.13 RFS attained with the expert WOD-E2E trajectories.